CONFIDENTIAL BFF/BFM/17181/ BFF1922 QUESTION 1 (25 marks) COMHI Motors manufactures high-performance battery electric vehicles. An engineer is on COMHI committee to evaluate bids for new generation coordinate-measuring machinery to be directly linked to the automated manufacturing of high-precision vehicle components. Three bids include the interest rates that vendors will charge on unpaid balances. To get a clear understanding of finance costs, COMHI management asked the engineer to determine the effective semiannual and annual interest rates for each bid. The bids are as follows: Bid 1 : Bid 2: Bid 3: 9% per year,compounded quarterly 3% per quarter, compunded quarterly 8.8% per year, compunded monthly Determine the effective rate for each bid on the basis of semiannual periods (10 marks) What are the effective annual rate? These are to be a part of the final bid selection (10 marks) (5 marks) a. b. c. Which bid has the lowest efective annual rate QUESTION 2 (25 marks) In-flight texting, phone, and internet connections provided at airline passenger seats are an expected service by many customers, M Airline knows it will have to replace 15,000 to
Answer #1

The effective rate can be calculated using the following formula

/0 Ef fectivem

a. Bid 1: Effective rate = 9%/2 = 4.5% semi annual

Bid 2: Effective rate = 3% per quarter

Semi annually effective rate = 6%

Bid 3: Effective rate = 8.8%/12 = 0.733% per month

Semi annually effective rate = 4.4% (Approximately)

b. Effective annual rate

Effective annual rate = (1+i)m - 1

Bid 1, EAR = (1+0.0225)4 - 1 = 1.093083 - 1 = 0.093083

Bid 1, effective annual rate = 9.31%

Bid 2, EAR = (1+ 0.03)4 - 1 = 0.1255

Bid 2, effective interest rate = 12.55%

Bid 3, EAR = (1+0.00733)12 - 1 = 1.09167 - 1 = 0.09167

Effective interest rate = 9.167%

c. Bid 3 has lowest effective annual rate.

Bid 1 effective annual rate = 9.31%

Bid 2 effective annual rate = 12.55%

Bid 3 effective annual rate = 9.17%

Please contact if having any query thank you.